OpenStudy (dinorap1):

Check my answers for my consumer math homework please! I will post the questions and answers below.

OpenStudy (dinorap1):

Here's the questions.

OpenStudy (anonymous):

The first one is correct: bi-weekly= paying every 2 weeks 2*133= 226 Let me look at the other ones..

OpenStudy (anonymous):

The second one seems correct, 499 divided by 6 equals $83

OpenStudy (dinorap1):

Thanks for your help so far. I'm pretty sure number one and two are right, since there's an example in the lesson that is exactly like those two questions. I'm just not sure about the last two.

OpenStudy (anonymous):

For the 3rd one I think it's the 2nd answer choice, you can have someone else check if you want. Onto the next one~

OpenStudy (dinorap1):

Hmmm what did you do for the math on that one?

OpenStudy (dinorap1):

Let me check

OpenStudy (anonymous):

And the 4th one seems correct too, 147 divided by 3(months)= 49-50 dollars You would have to pay 49 to 50 dollars monthly for car insurance ^_^

OpenStudy (dinorap1):

Here's what I did: 1850+113.5+113.5+98.6+100+398.45+83.33+80.16+80.16=2906.9

OpenStudy (dinorap1):

That's what I did for number 3. I think you must've forgotten one of the bi-weekly things when you were adding

OpenStudy (dinorap1):

Okay thank you for all your help! I will see what I get